OpenStudy (anonymous):

designing a circular carpet for his living room. He knows that half way across the carpet is 12 feet. He wants to put a rope border around the outside of the carpet. How many feet of rope would he have to order? so I'm confused to as if i use 24 instead of 12 and do i divide or multiply that with what to find the outside?

OpenStudy (anonymous):

diameter x pi?

jimthompson5910 (jim_thompson5910):

The radius is 12 feet. So the diameter is 12*2 = 24 feet The circumference is C = pi*d ... where d is the diameter which means C = pi*d C = pi*24 C = 24pi C = 24(3.14159) C = 75.39816 So he needs about 75.39816 feet of rope.

OpenStudy (anonymous):

whao i was alot off thank you

jimthompson5910 (jim_thompson5910):

You're welcome